OpenStudy (anonymous):

Find the derivative y' of the function: y=(x/(x^2+1))^4

OpenStudy (bahrom7893):

oh boy chain rule, quotient rule, power rule..

OpenStudy (anonymous):

is this implicit differentiation?

OpenStudy (bahrom7893):

no, you just have y on its own on the left. No need for that.

OpenStudy (mertsj):

\[y=\frac{x^4}{(x^2+1)^4}\] \[y'=\frac{(x^2+1)^4(4x^3)-(x^4)(4(x^2+1)^3(2x)}{(x^2+1)^8}\]

OpenStudy (mertsj):

You probably want to clean that up a little.

OpenStudy (anonymous):

wow it's that long...

OpenStudy (anonymous):

it will clean up a lot!

OpenStudy (mertsj):

No doubt.

OpenStudy (anonymous):

you could also use \[f'(x)=4\left(\frac{x}{x^2+1}\right)^3\times \frac{d}{dx}\left(\frac{x}{x^2+1}\right)\] but there is no avoiding the quotient rule somewhere

OpenStudy (anonymous):

just might be neater in the end use the chain rule. second term is \[\frac{-x^2+1}{(x^2+1)^2}\] so maybe less algebra at the end